OpenStudy (anonymous):

Find the distance between this (10,20) (13,16) Show your work. please its Geometry Btw.

OpenStudy (bahrom7893):

d^2 = (16-20)^2 + (13-10)^2 d^2 = 16 + 9 = 25 d = 5

OpenStudy (anonymous):

5

OpenStudy (anonymous):

distance formula

OpenStudy (anonymous):

Distance formula: d = √[ ( x - x')² + y - y')² ] = 5